a block of mass m1 fastened to a spring and d block of mass m2is placed against it on an inclined plane.the blocks are pushed to a dist of (2/k)(m1+m2)gsin@ against d spring and releasd . . find the position wher the 2 blocks separate?????

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Sneha,

ANS:-Let the velocity of the masses when the spring is again totally relaxesd is v, Then from the energy eq we get

1/2kd²=1/2(M1+M2)v²+(M1+M2)gdsinθ

Putting d=2/k(M1+M2)gsinθ & Solving we get v=0.

Hence the velocity is o, So, it will again compress the spring and let the comp is x then from energy eq we get,

(M1+M2)gsinθ x=1/2kx²

So again x=2(M1+M2)gsinθ/k

So, the two mass will execute Periodic motion with amplitude (M1+M2)gsinθ/k

THUS THEY WILL NEVER SEPARATE OUT

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sneha!!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR